# Quantum Dynamic Programming

### v1

@@ Abstract
can be achieved with a circuit of depth $\mathcal{O}(N^{2}\epsilon^{-1})$. Here, $\rho_{N}$ is the exact solution to the recursion.
The full, technical version of Theorem 1 can be found in Supplemental Materials Section III, as Theorem 9.
@@
Exponential width increase mitigation.—Realistically, quantum devices can neither execute exponentially many sequential gates nor operate on exponentially many qubits simultaneously. Therefore, it becomes inevitable to consider a hybrid approach, where QDP is initiated after several rounds of unfolding recursions. This hybrid approach strategically allocates the exponential factor between circuit depth and width, without overburdening either. One can envisage a scheme where $N_{1}$ recursions unfold, using a circuit of depth $e^{\mathcal{O}(N_{1})}$, nearly reaching the device’s limit. These unfolding steps, conducted in parallel to exploit maximum width, yield $M$ copies of intermediate states $\rho_{N_{1}}$. Subsequently, QDP takes over, executing an additional $N_{2}$ recursions to attain the final state $\sigma_{N_{1}+N_{2}}$. This strategy maximizes the device’s capacity, which would have otherwise produced either $\rho_{N_{1}}$ without QDP or $\sigma_{N_{2}}$ without unfolding.
Another advantage of adopting a hybrid strategy arises from inherent implementation imperfection that QDP entails. As will be demonstrated in the Grover search example, this might limit the initial states to those already possessing certain favorable properties. In the hybrid approach, we effectively steer the ineligible initial state to meet the requirements imposed by QDP through initial rounds of unfolding recursions.
Example 1: Grover search.—Theorem 1, when applied to the fixed-point Grover search, indicates a circuit depth scaling of $\mathcal{O}(N^{2}\epsilon^{-1})$. However, a more refined analysis reveals a better scaling of $\mathcal{O}(N\epsilon^{-1})$. Specifically, we analyze a QDP implementation of the nested fixed-point Grover search [11], a recursive variant of Grover’s fixed-point algorithm [15], characterized by Eq. (5). This operation transforms the input state into a state closer to the target:
$\left\lVert\tau-\hat{\Gamma}^{(\psi)}_{L}\psi\left(\hat{\Gamma}^{(\psi)}_{L}\right)^{\dagger}\right\rVert_{1}<\bigg{\lVert}\tau-\psi\bigg{\rVert}_{1}\ .$ (13)
For a single recursion step, as analyzed in Ref. [11], the output state with a distance $\delta$ to the target state is achieved by $L=\mathcal{O}(\log(1/\delta))$. The circuit depth $\mathcal{O}(L)=\mathcal{O}(\log(1/\delta))$ matches that of the optimal Grover search [19, 20].
Ref. [11] additionally shows that the same distance $\delta$ can be achieved by executing $N$ recursions with shorter length $L=\mathcal{O}((\log(1/\delta))^{\frac{1}{N}})$, following the recurrence relation $|{\psi_{n+1}}\rangle=\hat{\Gamma}^{(\psi_{n})}_{L}|{\psi_{n}}\rangle$ as discussed earlier. The recursion is implemented through unfolding, and the total circuit depth for $N$ iterations scales exponentially as $\mathcal{O}((2L+1)^{N})=\mathcal{O}(\log(1/\delta))$, asymptotically equivalent to the non-recursive version of the algorithm; refer to Supplemental Materials Sec. IV.1 for a self-contained discussion. In contrast, our QDP implementation reduces the circuit depth to $\mathcal{O}(N)$, implying an exponential runtime speedup for the implementation of Grover search.
Theorem 2 (Efficient QDP Grover search, non-technical).
Suppose that the initial state $\psi_{0}$ is sufficiently close to the target state $\tau$ for the nested fixed-point Grover search. By implementing $N$ recursion steps via QDP, one can prepare the final state that is $\epsilon$-close to that of $N$ exact recursion steps. This implementation uses a circuit of depth $\mathcal{O}(N\epsilon^{-1})$ and $e^{\mathcal{O}(N)}$ copies of the initial state $\psi_{0}$.
The technical version of Theorem 2 is Theorem 21 in Supplemental Materials Sec. IV, and we provide the detailed analysis there. The circuit depth scaling is better than in Theorem 1 thanks to an additional protocol based on Ref. [21] which probabilistically reduces the mixedness of a state, while preserving its eigenvectors.
The requirement of small $\lVert\tau-\psi_{0}\rVert_{1}$ stems from approximate memory-calls in QDP: when the initial overlap is too small, the contraction in the distance to the target state may be insufficient to offset the implementation error, i.e. the algorithm lacks fast spectral convergence. If the given initial state does not meet this criterion, a hybrid approach should be employed instead.
Example 2: Oblivious Schmidt decomposition.—For any bi-partite pure state $|{\psi_{0}}\rangle\in\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, there exist local unitaries $\hat{V}_{A}$ and $\hat{V}_{B}$ that simplify the Schmidt decomposition to be in the computational basis, i.e. $\hat{V}_{A}\otimes\hat{V}_{B}|{\psi_{0}}\rangle=\sum_{k=1}^{D}\sqrt{\lambda_{k}}|{k}\rangle\otimes|{k}\rangle$, where $\{\lambda_{k}\}_{k}$ denotes the Schmidt spectrum [22, 23]. This decomposition is of paramount importance for quantum information theory but of limited use as a practical technique. The main obstacle is that $\hat{V}_{A}$ and $\hat{V}_{B}$ depend on the state $|{\psi_{0}}\rangle$; thus, it seemed necessary to learn $|{\psi_{0}}\rangle$ first, classically compile the unitaries, and implement them in physical devices. Instead, QDP allows the oblivious execution of the Schmidt decomposition, without requiring a classical description of $|{\psi_{0}}\rangle$, by applying double-bracket iterations [12] to the reduced state of $|{\psi_{0}}\rangle$.
We describe the gist of the algorithm here, and expound further in Supplemental Materials Section V. Let $\hat{D}$ be a diagonal operator on subsystem $A$. The oblivious Schmidt decomposition recursion is defined as $|{\psi_{n+1}}\rangle=\hat{U}^{(\psi_{n})}|{\psi_{n}}\rangle$, where
$\hat{U}^{(\psi_{n})}=e^{s[\hat{D},\rho_{n}^{(A)}]}\otimes\mathmybb{1}_{B}\ ,$ (14)
and $\rho_{n}^{(A)}=\operatorname{Tr}_{B}[|{\psi_{n}}\rangle\!\langle{\psi_{n}}|]$ is the reduced density matrix after $n$ steps. In other words, it is a single memory-call type recursion with $\hat{\mathcal{N}}(\rho)=-is[\hat{D},\rho]$ that is Hermitian-preserving. If $\hat{D}$ has a non-degenerate spectrum and $s$ is small [24], the reduced state of the fixed-point is a diagonal state isospectral to $\rho_{0}^{(A)}$. Under the same condition, exponential convergence can be proven, which is conceptually similar to fast spectral convergence and is a feature of gradient flows in general [25]. One can adapt $\hat{D}_{n}$ in every step variationally, and then in practice, a non-degenerate spectrum is not required [12].
In QDP implementation, Eq. (14) is generated by coupling many copies of $\rho_{n}^{(A)}$, prepared independently each time starting from $\rho_{0}^{(A)}=\operatorname{Tr}_{B}[|{\psi_{0}}\rangle\!\langle{\psi_{0}}|]$. After $N$ iterations, at which $\rho_{N}^{(A)}$ is sufficiently diagonal, the Schmidt spectrum can be extracted via measurements in the computational basis.
@@
Obtaining the Schmidt spectrum of a density matrix can be done also by the ‘replica’ method [26], i.e. by measuring $\operatorname{Tr}(C_{A}\psi_{AB}^{\otimes k})=\operatorname{Tr}\rho_{A}^{k}$, where $C_{A}|{i_{1},i_{2},\dots,i_{k}}\rangle_{A}=|{i_{k},i_{1},i_{2},\dots,i_{k-1}}\rangle_{A}$ is the cycle permutation of $k$ copies of the subsystem $A$ and performing classical data analysis of $\operatorname{Tr}\rho_{A}^{k}$ for $k=2,\dots,d$. Besides requiring exponentially many swap operations for $k=d$, the replica approach is incoherent. The major strength of oblivious Schmidt decomposition is to provide coherently the bipartite quantum state $\sum_{k}\sqrt{\lambda_{k}}|{k}\rangle\otimes|{k}\rangle$ in the computational basis. This is useful, e.g. for entanglement distillation of an unknown state $\psi$, which contrasts to the standard setting [27] where the information about the initial state is necessary for compiling appropriate local operations. Ref. [28] showed that the same asymptotic rate (characterized by entanglement entropy) can be achieved for the oblivious setting. Oblivious Schmidt decomposition may not realize such an optimal rate but provides a constructive way of oblivious entanglement distillation, which comes with an explicit circuit construction.
@@
Conclusion and outlook.— We introduced the general framework of quantum dynamic programming that, similarly to the classical version, leverages placeholder memory to achieve runtime reductions. We found that it is capable of implementing a broad class of quantum recursions involving memory-calls. Whenever fixed-points of recursions are sufficiently well-behaved, we showed that QDP yields accurate results with polynomial depth. This represents an exponential speed-up over the previously proposed zero-memory implementations by ‘unfolding’ [11, 12]. Our approach introduces a new possibility of balancing circuit depth and width, facilitating the implementation of quantum recursions in real devices. QDP casts quantum recursions into circuits with many parallel but few sequential operations and renders them suitable as a central use-case for distributed quantum computing [7, 8, 9].
@@
For QDP applications involving only a marginal amount of quantum coherence, an approximate cloning heuristic [29] could be used: these cases would potentially enjoy an exponential speed-up without inflating the circuit width. Nevertheless, quantitative bounds relating the coherence of input quantum states to cloning fidelity are currently unknown [30]. It is also unclear how far the presence of coherence could be viewed as a signature of quantum simulation hardness, generalizing to quantum recursions the role of entanglement in tensor network simulations [31]. Thorough investigations on the role of coherence in QDP will make an interesting future direction.
@@
We do not know of quantum algorithms for which QDP facilitates a rigorous computational advantage over classical algorithms. These may appear in settings robust against imprecision of unitary implementations, similar to those present in diagonalizing double-bracket iterations [24, 25, 12]. Indeed, QDP allows us to add oblivious Schmidt decomposition to the available quantum algorithmic toolkit. It is our hope that once it will be feasible to experimentally implement memory-usage queries with high fidelity, then oblivious Schmidt decomposition and QDP in general will facilitate practical state preparations that will advance our knowledge about quantum properties of materials, e.g. magnets or superconductors.
@@
Acknowledgements.—JS, MG, and NN are supported by the start-up grant for Nanyang Assistant Professorship of Nanyang Technological University, Singapore, awarded to Nelly H. Y. Ng. MG additionally acknowledges support through the Presidential Postdoctoral Fellowship of the Nanyang Technological University. RT acknowledges the support of JSPS KAKENHI Grant Number JP23K19028, JST, CREST Grant Number JPMJCR23I3, Japan, and the Lee Kuan Yew Postdoctoral Fellowship of Nanyang Technological University Singapore.
@@ I Preliminaries and distance measures
$\frac{1}{2}\lVert\hat{A}-\hat{B}\rVert_{1}\coloneq\frac{1}{2}\operatorname{Tr}\left[\sqrt{(\hat{A}-\hat{B})^{\dagger}(\hat{A}-\hat{B})}\right]\ .$ (15)
@@
$\frac{1}{2}\lVert\hat{A}-\hat{B}\rVert_{2}\coloneq\frac{1}{2}\sqrt{\operatorname{Tr}\left[(\hat{A}-\hat{B})^{\dagger}(\hat{A}-\hat{B})\right]}\ .$ (16)
@@
$\frac{1}{2}\|\Phi_{1}-\Phi_{2}\|_{\rm Tr}\coloneq\frac{1}{2}\max_{\varrho\in\mathcal{S}(\mathcal{H})}\left\lVert\Phi_{1}(\varrho)-\Phi_{2}(\varrho)\right\rVert_{1}\ .$ (17)
@@
$\frac{1}{2}\|\Phi_{1}-\Phi_{2}\|_{\diamond}\coloneq\frac{1}{2}\max_{\varrho\in\mathcal{S}(\mathcal{H}\otimes\mathcal{H})}\left\lVert(\Phi_{1}\otimes\operatorname{id})(\varrho)-(\Phi_{2}\otimes\operatorname{id})(\varrho)\right\rVert_{1}\ .$ (18)

### v2

@@ Abstract
using a circuit of depth $\mathcal{O}(N\epsilon^{-1})$ and width $\epsilon^{-N}e^{\mathcal{O}(N)}$. Here, $|{\psi_{N}}\rangle$ is the exact solution to the recursion.
The technical version of this theorem and its proof is given as Theorem 12 in [27]. We prove this theorem by addressing two different types of error. In particular, we note that unitary errors are efficiently suppressed by fast spectral convergence, while non-unitary errors destabilize the convergence and need to be mitigated explicitly. We employ an additional protocol [30, 31] to reduce such errors, and show in [27] Sec. III.2 that this protocol incurs only a slightly faster exponential growth in circuit width. Furthermore, a similar theorem can be developed for mixed state extensions, which we display in End Matters.
Implications of the depth-width trade-off: distributed quantum computing and hybrid strategy.—Theorem 1 establishes the desired circuit depth-width trade-off. At its core, this trade-off is valuable because it exponentially reduces computational time, which can turn a practically impossible task into a feasible one. In addition, the capacity to realize a shorter depth circuit comes with potential advantages when combined with quantum error mitigation. In particular, it was shown that error mitigation generally requires exponential resource overhead with circuit depth [32, 33, 34]. Although there are observations indicating that error mitigation could require exponential cost also with circuit width [35, 36], a rigorous characterization applicable to general classes of circuits and error mitigation methods has not been established. In this sense, there is still room for some error mitigation techniques to run efficiently in large-width short-depth circuits of practical interest. Beyond this immediate advantage, we focus on how QDP facilitates implementation.
One simple measure is the quantum circuit size, defined as the product of the maximum circuit depth and width [37, 38]. We illustrate how this measure benefits from the trade-off. Suppose that in each recursion step, $L$ non-covariant memory-calls are made, each approximated by $k$ covariant ones. The unfolding circuit size therefore scales as $(2kL+1)^{N}$ for $N$ recursion steps. Meanwhile, memory-usage queries directly implement even non-covariant memory-calls, but require $k^{\prime}$ memory-usage queries to manage implementation error, i.e. $M=k^{\prime}L$ memory-usage queries per recursion step. In this scenario, the QDP circuit width scales as $(M+1)^{N}=(k^{\prime}L+1)^{N}$, which dominates the total circuit size for large $N$. Whenever $L$ is large and $k^{\prime}<2k$, QDP approximately achieves a $(2k/k^{\prime})^{N}$-fold improvement in circuit size compared to unfolding. For instance, in the most error-tolerant case of $k=k^{\prime}=1$, QDP significantly reduces circuit size.
Beyond circuit size, QDP offers additional practical advantages: its local modularity is particularly well-suited for distributed quantum computing [39, 40, 41]. By localizing and decoupling circuits [42], QDP allows copies of $|{\psi_{n}}\rangle$ to be prepared independently before being injected into recursion steps $\hat{U}^{[\psi_{n}]}$. As the recursion progresses, the circuit width decreases exponentially as most copies are consumed as memory and traced out during memory-usage queries. This enables greater flexibility in parallelization, since processors with shorter coherence times can be allocated for preparing $|{\psi_{n}}\rangle$ with smaller $n$, optimizing the use of available quantum devices.
@@
The flexibility of QDP is reinforced by a hybrid strategy. Realistically, quantum devices can neither execute exponentially many sequential gates nor operate on exponentially many qubits simultaneously. A practical approach is to initiate QDP after several rounds of unfolding, distributing the exponential factor between circuit depth and width to prevent either from becoming prohibitive. For example, one could begin with $N_{1}$ unfolding steps, using a circuit of depth $e^{\mathcal{O}(N_{1})}$, nearing the device’s depth limit. These unfolding steps are performed in parallel, maximizing circuit width to produce $M$ copies of intermediate states $|{\psi_{N_{1}}}\rangle$. For many quantum recursions, this starting strategy also conveniently steers the system into a regime where the recursion starts to exhibit fast spectral convergence. This allows QDP to subsequently take over, executing an additional $N_{2}$ recursions to attain the final state $\sigma_{N_{1}+N_{2}}$. This hybrid strategy fully utilizes the device capacity, which is otherwise limited to producing either $|{\psi_{N_{1}}}\rangle$ without QDP or $\sigma_{N_{2}}$ without unfolding.
Example 1: Grover search.—The nested fixed-point Grover search [5] has a recursion unitary defined in Eq. (3), where $N$ recursion steps of this algorithm prepare $|{\psi_{N}}\rangle$ such that
$\frac{1}{2}\lVert|{\psi_{N}}\rangle\!\langle{\psi_{N}}|-|{\tau}\rangle\!\langle{\tau}|\rVert_{1}\simeq e^{-\lvert\langle{\psi_{0}}|{\tau}\rangle\rvert(2L+1)^{N}}\ .$ (9)
If the recursion is implemented through unfolding, the total circuit depth scales exponentially as $(2L+1)^{N}$. This matches the performance of the optimal Grover search [25, 43] with the same circuit depth; refer to [27] Sec. IV.1 for a self-contained discussion. In contrast, the QDP implementation reduces the circuit depth to $\mathcal{O}(N)$, implying an exponential runtime speedup, thanks to Eq. (9) exhibiting fast spectral convergence. This speedup comes with the requirement of exponential copies of the initial state $|{\psi_{0}}\rangle$. Unfortunately, it is known that [16] there is no quantum advantage in sample-based Grover search algorithms, which include the QDP implementation. Nevertheless, this example serves as a valuable demonstration of QDP’s applicability for a historically significant quantum algorithm.
Example 2: Quantum imaginary-time evolution.—We consider quantum imaginary-time evolution in the sample-based Hamiltonian simulation scenario [16], which aims at achieving the task of ground state preparation for quantum many-body systems. Suppose that copies of $\rho\propto\hat{H}$ are given, where $\hat{H}$ is the Hamiltonian with a zero ground state energy. We also prepare copies of some pure state $|{\psi_{0}}\rangle$ having non-zero overlap with the ground state. Using them, recursions of the form
$|{\psi_{n}}\rangle\mapsto|{\psi_{n+1}}\rangle=e^{s[|{\psi_{n}}\rangle\!\langle{\psi_{n}}|,\hat{H}]}|{\psi_{n}}\rangle,$ (10)
with some duration $s$ can be implemented oblivious to both $\hat{H}$ and $|{\psi_{n}}\rangle$; see [27] Secs. V.3 and V.4. This unitary has been shown [9] to be a discretized version of the quantum imaginary-time evolution [44] and yield a sequence of states $\{|{\psi_{n}}\rangle\}$ whose infidelity to the ground state decays exponentially as $n$ grows, i.e. with fast spectral convergence. An alternative QDP implementation of the same task with the usual Hamiltonian simulation, instead of access to copies of $\rho\propto\hat{H}$, is studied in Ref. [9].
Example 3: Oblivious Schmidt decomposition.—For pure bipartite states, the Schmidt decomposition [45, 46] fully characterizes their entanglement. Thus we may envision a protocol that unitarily transforms any unknown $|{\psi_{0}}\rangle\in\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ into a form where its Schmidt basis aligns with the computational basis: $\hat{V}_{A}\otimes\hat{V}_{B}|{\psi_{0}}\rangle=\sum_{k=1}^{D}\sqrt{\lambda_{k}}|{k}\rangle\otimes|{k}\rangle$ with $\{\lambda_{k}\}_{k}$ denoting the Schmidt spectrum. Traditionally, prior knowledge of $|{\psi_{0}}\rangle$ is required to construct $\hat{V}_{A}$ and $\hat{V}_{B}$. However, QDP enables this task while completely circumventing such a need. Given that this is a novel algorithm, we outline the key ideas here, and provide a more detailed exposition in [27] Sec. V.5. Let $\hat{D}$ be a non-degenerate diagonal operator on subsystem $A$. The recursion unitary is defined as
$\hat{U}^{[\hat{\mathcal{N}},\psi_{n}]}=e^{s\left[\hat{D},\operatorname{Tr}_{B}[|{\psi_{n}}\rangle\!\langle{\psi_{n}}|]\right]}\otimes\mathmybb{1}_{B}\ .$ (11)
This is a single memory-call type recursion with $\hat{\mathcal{N}}(\rho_{AB})=-is[\hat{D},\operatorname{Tr}_{B}[\rho_{AB}]]\otimes\mathmybb{1}_{B}$ that is Hermitian-preserving; hence, with copies of $|{\psi_{n}}\rangle$, the recursion step $|{\psi_{n}}\rangle\mapsto|{\psi_{n+1}}\rangle=\hat{U}^{[\psi_{n}]}|{\psi_{n}}\rangle$ can be approximated. When $s$ is small, $\{|{\psi_{n}}\rangle\}$ exponentially converges to $\hat{V}_{A}\otimes\mathmybb{1}_{B}|{\psi_{0}}\rangle$ [47, 48] allowing Theorem 1 to apply. A similar procedure implements $\hat{V}_{B}$, completing the oblivious Schmidt decomposition.
@@
Let us discuss the implications of oblivious Schmidt decomposition in the context of quantum information processing. The replica method [49] extracts classical information $\lambda_{k}$ but requires exponentially many swap operations. Similarly, we expect that oblivious Schmidt decomposition may also require a long runtime to converge. However, it not only enables sampling from the Schmidt spectrum but also provides the bipartite quantum state $\sum_{k}\sqrt{\lambda_{k}}|{k}\rangle\otimes|{k}\rangle$ coherently in the computational basis. This is useful, e.g. for entanglement distillation of an unknown state $\psi$, contrasting with standard settings [50] that require the knowledge about the initial state for compiling local operations. While oblivious Schmidt decomposition via QDP may not attain the optimal asymptotic rates as derived from Ref. [51], it provides a constructive approach to oblivious entanglement distillation with explicit circuit implementations.
@@
Conclusion and outlook.—We demonstrated that QDP enables the efficient implementation of a broad class of quantum recursions involving memory-calls. For well-behaved recursions, QDP yields accurate results with polynomial depth—an exponential speed-up over the previously proposed zero-memory implementations based on unfolding [5, 6]. Additionally, QDP enables hybrid strategies of balancing circuit depth and width, making quantum recursions more feasible on real devices. By structuring quantum recursions into circuits with parallel rather than sequential operations, QDP renders them suitable as a central use-case for distributed quantum computing [39, 40, 41].
@@
The general QDP proposal assumes the preparation of many independent copies of the same quantum state, analogous to sample-based algorithms [13, 15, 16, 52, 17, 22, 23], where copies of the quantum state replace oracle access to unitary operators. For QDP applications involving only a marginal amount of quantum coherence, approximate cloning heuristics [21] could potentially provide exponential speed-ups without inflating the circuit width. Nevertheless, quantitative bounds relating the coherence of input quantum states to cloning fidelity are currently unknown [53]. It is also unclear whether coherence could signal simulation hardness for quantum recursions, akin to how entanglement affects tensor network simulations [54]. Investigating the role of coherence in QDP presents an intriguing avenue for future research.
@@
Currently, we are unaware of quantum algorithms where QDP ensures a clear computational advantage over classical methods. These may appear in settings robust against imperfect unitary implementations, similar to those present in diagonalizing double-bracket iterations [47, 48, 6]. Indeed, QDP allows us to add oblivious Schmidt decomposition to the quantum algorithmic toolkit. We hope that once it will be feasible to experimentally implement memory-usage queries with high fidelity, oblivious Schmidt decomposition and QDP in general will facilitate practical state preparations that will advance our knowledge of quantum properties in materials, e.g. magnets or superconductors.
@@
Acknowledgements.—We thank Debbie Leung for drawing our attention to noise reduction algorithms applicable to higher dimensions. JS, MG, and NN are supported by the start-up grant for Nanyang Assistant Professorship of Nanyang Technological University, Singapore. MG additionally acknowledges support through the Presidential Postdoctoral Fellowship of the Nanyang Technological University. RT acknowledges the support of JSPS KAKENHI Grant Number JP23K19028, JST, CREST Grant Number JPMJCR23I3, Japan, and the Lee Kuan Yew Postdoctoral Fellowship of Nanyang Technological University Singapore.
@@ I Preliminaries and distance measures
$\frac{1}{2}\lVert\hat{A}-\hat{B}\rVert_{1}\coloneq\frac{1}{2}\operatorname{Tr}\left[\sqrt{(\hat{A}-\hat{B})^{\dagger}(\hat{A}-\hat{B})}\right]\ .$ (S1)
@@
$\frac{1}{2}\lVert\hat{A}-\hat{B}\rVert_{2}\coloneq\frac{1}{2}\sqrt{\operatorname{Tr}\left[(\hat{A}-\hat{B})^{\dagger}(\hat{A}-\hat{B})\right]}\ .$ (S2)
@@
$\frac{1}{2}\|\Phi_{1}-\Phi_{2}\|_{\rm Tr}\coloneq\frac{1}{2}\max_{\varrho\in\mathcal{S}(\mathcal{H})}\left\lVert\Phi_{1}(\varrho)-\Phi_{2}(\varrho)\right\rVert_{1}\ .$ (S3)
@@
$\frac{1}{2}\|\Phi_{1}-\Phi_{2}\|_{\diamond}\coloneq\frac{1}{2}\max_{\varrho\in\mathcal{S}(\mathcal{H}\otimes\mathcal{H})}\left\lVert(\Phi_{1}\otimes\operatorname{id})(\varrho)-(\Phi_{2}\otimes\operatorname{id})(\varrho)\right\rVert_{1}\ .$ (S4)